# Geometric Signatures of Compositionality Across a Language Model’s Lifetime

### v2

@@ Appendix I Additional Results: Per-Layer Correlation with Kolmogorov Complexity
Figure I.1: Spearman correlations between per-layer dimensionality and estimated Kolmogorov complexity. The Spearman correlation between the gzipped dataset size (KB) and representational dimensionality per layer, is shown for all tested model sizes. Generally across models, per-layer $I_{d}$ is not correlated to the estimated Kolmogorov complexity, or formal compositionality, of datasets. Per-layer linear $d$ is consistently highly positively correlated to the estimated Kolmogorov complexity, except one outlier (160m).
@@
Figure I.2: Spearman correlation of layerwise dimensionality and Kolmogorov complexity over training. The Spearman correlations between $I_{d}$ (left) and gzip and $d$ (right) and gzip are plotted for three models, 410m, 1.4b, 6.9b (top to bottom) over training time (x axis), where correlation is computed across the controlled corpora. Each vertical set of points denotes the layer distribution of Spearman correlations at a single timestep; each point is one layer’s Spearman correlation, colored green if statistically significant and gray otherwise.
@@
There was one significant outlier, 160m, in our analysis correlating layerwise dimensionality to gzip (Kolmogorov complexity), see Figure I.1 and Table 1. While other models consistently demonstrate a positive Spearman correlation between $d$ and gzip across layers, 160m (and to a smaller extent, 70m) deviates from this pattern. The reason 160m displays a negative correlation is due to its behavior on shuffled corpora, see the third column in Figure G.1: for intermediate layers, PCA with a variance threshold of 0.99 yields fewer than 50 PCs. We found that this was due to the existence of so-called “rogue dimensions" (Timkey & van Schijndel, 2021; Machina & Mercer, 2024; Rudman et al., 2023), where very few dimensions have outsized norms. Outlier dimensions have been found, via mechanistic interpretability analyses, to serve as a “sink" for uncertainty, and are associated to very frequent tokens in the training data (Puccetti et al., 2022). See Rudman et al. (2023) for exact activation profiles for the last-token embeddings in Pythia 70m and 160m. While increasing the variance threshold to $0.999$ reduced the effect of rogue dimensions on PCA dimensionality estimation, we decided to keep the threshold at $0.99$ for consistent comparison to other models.
Coding of formal complexity over training
@@
    PCA encodes formal complexity (seen by earlier dots close to $\rho=1.0$) as an inductive bias of the model architecture. The high correlation for most layers may be unlearned during intermediate checkpoints of model training, seen by the “dip" in gray dots around steps $10^{2}\sim 10^{3}$, but is regained by the end of training for all model sizes. This indicates that encoding formal complexity at the end of training is a learned behavior.
@@
    For $I_{d}$, the phase transition noted in Section 4.2 is also present at slightly before $t=10^{3}$; this is seen by layerwise correlations in Figure I.2a coalescing to around $\rho=0.5$, and then redistributing. The layers that best encode formal complexity for TwoNN at the end of trianing correspond to model-initial and model-final layers, see Figure I.1 top.

### v3

@@ I.1 Correlation between Kolmogorov Complexity and Feature Complexity is Robust to Sequence Length
Table I.1: Spearman correlations between dimensionality and estimated Kolmogorov complexity, varying sequence length. The Spearman correlation $\rho$ between the gzipped dataset size (KB) and representational dimensionality (rows), averaged over layers, is shown for all tested Pythia model sizes (model name omitted for readability). Values marked with a * are significant with a p-value threshold of 0.05. Values marked with † are significant with a p-value threshold of 0.1. Across models, average-layer $I_{d}$ is not correlated to the estimated KC, or formal compositionality, of datasets. Average-layer linear $d$ is consistently highly positively correlated to the estimated KC. Length $l=5$ is grayed out as, due to the sequence length being too short, it was not possible to varying the coupling factor $k$; here, the only comparison is between coherent and shuffled ($n=2$).
sequence length (words)
5 8 11 15 17
14m $I_{d}$ 1.00 0.89 0.87 0.87 -0.10
$d$ 1.00 0.89 0.87 0.87 0.81
70m $I_{d}$ 1.00 0.40 0.43 0.26 -0.10
$d$ 1.00 1.00 1.00 0.98 0.98
160m $I_{d}$1.00 0.00 0.19 0.00 -0.21
$d$1.00 -0.60 -0.52 -0.62 -0.62
410m $I_{d}$1.00 0.40 0.40 0.26 0.14
$d$ 1.00 1.00 0.98 1.00 1.00
1.4b $I_{d}$1.00 0.40 0.40 0.43 0.14
$d$ 1.00 1.00 0.98 1.00 1.00
6.9b $I_{d}$ 1.00 0.40 0.40 0.36 0.48
$d$ 1.00 1.00 0.98 1.00 0.98
12b $I_{d}$ 1.00 0.40 0.40 0.43 0.00
$d$ 1.00 1.00 0.98 1.00 1.00
Llama-8b $I_{d}$ 1.00 0.40 0.19 0.00 -0.02
$d$ 1.00 1.00 0.98 1.00 0.93
Mistral-7b $I_{d}$ 1.00 0.40 0.40 0.00 0.29
$d$ 1.00 1.00 0.98 1.00 0.90
Figure I.1: Average layerwise dimensionality vs. Estimated Kolmogorov Complexity (gzip) for Pythia 410m, 1.4b, and 6.9b, aggregated for all grammars. For all models, PCA $d$ highly correlates to gzip (estimated KC), with Spearman $\rho\geq 0.9$** for all models. TwoNN $I_{d}$ correlates more weakly, $\rho\in[0.5,0.6]$* for all models. Linear $d$ and nonlinear $I_{d}$ differentially encode shuffled data complexity (orange dots) compared to coherent data complexity (blue dots); where shuffled data display higher $d$ and lower $I_{d}$. (**) Significant at $\alpha=0.001$, (*) $\alpha=0.01$.
I.3 Per-layer Correlation with Kolmogorov Complexity
Figure I.2: Spearman correlations between per-layer dimensionality and estimated Kolmogorov complexity, Pythia models. The Spearman correlation between the gzipped dataset size (KB) and representational dimensionality per layer, is shown for all tested model sizes for the longest sequence length ($l=17$). Generally across models, per-layer $I_{d}$ is not correlated to the estimated KC, or formal compositionality, of datasets. Per-layer linear $d$ is consistently highly positively correlated to the estimated Kolmogorov complexity, except one outlier (160m).
Figure I.3: Spearman correlations between per-layer dimensionality and estimated Kolmogorov complexity, Llama-3-8B and Mistral-7B. The Spearman correlation between the gzipped dataset size (KB) and representational dimensionality per layer, is shown for Llama (left) and Mistral (right). Consistently across models, mirroring trends for Pythia, per-layer $I_{d}$ is not correlated to the estimated KC, or formal compositionality, of datasets. Per-layer linear $d$ is consistently highly positively correlated to the estimated KC.
Figure I.4: Per-layer Spearman $\rho$ between feature complexity and KC for all models, across all tested datasets. The layerwise Spearman $\rho$ between KC, measured with gzip, and feature complexity, measured with TwoNN $I_{d}$ (blue) and PCA $d$ (orange), is shown for each model. Each datapoint in each distribution corresponds to one (model, dataset, layer) triple. Generally across models, except for the outlier Pythia-160m, the layerwise correlation between $I_{d}$ and KC is low, while the correlation to $d$ is high and close to $1.0$ for the vast majority of layers, datasets, and models (orange distributions near $1.0$). This shows that, with high generality across models and datasets, the vast majority of layers encode KC in linear effective $d$, not in the intrinsic dimension $I_{d}$. Trends are especially robust after a certain model size ($\geq$410m).
@@ I.3 Per-layer Correlation with Kolmogorov Complexity
Figure I.5: Spearman correlation of layerwise dimensionality and Kolmogorov complexity over training. The Spearman correlations between $I_{d}$ (left) and gzip, and $d$ (right) and gzip are plotted for three models: 410m, 1.4b, and 6.9b (top to bottom) over training time (x-axis). Correlations are computed across the controlled corpora. Each vertical set of points denotes the layer distribution of Spearman correlations at a single timestep; each point is one layer’s Spearman correlation, colored green if statistically significant and gray otherwise.
Linear effective $d$ encodes KC robustly across models and datasets
Figure I.4 shows, for each model, the Spearman correlation between layer dimension and KC (gzip). Orange boxplots correspond to $d$, and blue boxplots to the $I_{d}$. Each datapoint in a boxplot reports the correlation for one (model, layer, sequence length) combination; the only factor of variation in each correlation is the $k$-coupling factor and whether the dataset is shuffled. With high generality across models and grammars, the linear effective $d$ is monotonic in KC, seen by the vast majority of layers (orange distributions) close to $\rho=1.0$ (y-axis). Meanwhile, the $I_{d}$ does not consistently encode superficial complexity, seen by the blue distributions landing about $0.0$.
@@
There was one significant outlier, 160m, in our analysis correlating layerwise dimensionality to gzip (KC), see Figures I.2 and I.4 and Table I.1. While other models consistently demonstrate a positive Spearman correlation between $d$ and gzip across layers, 160m (and to a smaller extent, 70m) deviates from this pattern. The reason 160m displays a negative correlation is due to its behavior on shuffled corpora, see the third column in Figure G.1: for intermediate layers, PCA with a variance threshold of 0.99 yields fewer than 50 PCs. We found that this was due to the existence of so-called “rogue dimensions" (Timkey and van Schijndel, 2021; Machina and Mercer, 2024; Rudman et al., 2023), where very few dimensions have outsized norms. Outlier dimensions have been found, via mechanistic interpretability analyses, to serve as a “sink" for uncertainty, and are associated to very frequent tokens in the training data (Puccetti et al., 2022). See Rudman et al. (2023) for exact activation profiles for the last-token embeddings in Pythia 70m and 160m. While increasing the variance threshold to $0.999$ reduced the effect of rogue dimensions on PCA dimensionality estimation, we decided to keep the threshold at $0.99$ for consistent comparison to other models.
Coding of superficial complexity over training
@@
    PCA encodes superficial complexity (seen by earlier dots close to $\rho=1.0$) as an inductive bias of the model architecture. The high correlation for most layers may be unlearned during intermediate checkpoints of model training, seen by the “dip" in gray dots around steps $10^{2}\sim 10^{3}$, but is regained by the end of training for all model sizes. This indicates that encoding superficial complexity at the end of training is a learned behavior.
@@
    For $I_{d}$, the phase transition noted in Section 4.2 is also present at slightly before $t=10^{3}$; this is seen by layerwise correlations in Figure I.5a coalescing to around $\rho=0.5$, and then redistributing. The layers that best encode superficial complexity for TwoNN at the end of trianing correspond to model-initial and model-final layers, see Figure I.2 top.
Appendix J Additional Results: Varying Sequence Length
On $d$ and $I_{d}$ computed for grammars of varying lengths, we found several behaviors that are emergent with sequence length:
  1. 1.
    Feature complexity increases with sequence length (Figure J.1). This is to be expected, as sequences that are longer contain more information that the LM needs to represent.
  2. 2.
    Shuffling feature collapse is an emergent property of sequence length. Figure J.2 shows the gap between coherent and shuffled $I_{d}$ grows as sentence length grows. Intuitively, this suggests that phrase-level feature complexity in short coherent sentences proxies that of a bag of words (shuffled).
We also find behaviors, corresponding to main article Section 4.3, that are robust to sequence length:
  1. 1.
    Feature complexity increases as $k$ decreases, no matter the sequence length (Figure J.3), showing that results presented in the main generalize to sequences of different length.
  2. 2.
    Upon shuffling, $I_{d}$ collapses to a low range, while $d$ increases, no matter the sequence length (Figure J.3).
[…]